90 men and 60 women were asked if they had watched the latest ‘Expendables’ movie. Altogether 3/5 of the people said Yes. 3/10 of the women said Yes. What percentage of the men said No?


1
Expert's answer
2021-10-28T07:03:04-0400

"\displaystyle
\text{The number of people that have watched expendables is }\\
\frac{3}{5} \cdot 150= 90\\
\text{The number of women that have watched the movie is }\\
\frac{3}{10}\cdot 60 = 18\\
\text{Therefore the number of men that have watched the movie is 90 - 18 = 72}\\
\text{The number of men that have not watched the movie is 90 -72 = 18}\\
\text{Hence the percentage of men that have not watched the movie is}\\
\frac{18}{150} \cdot 100=12\%"